OpenStudy (anonymous):

A one-ton ice sculpture is melting so that it loses one-fifth of its weight per hour. How much of the sculpture will be left after five hours ? write it in pounds

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

First convert 1 ton into pounds!

OpenStudy (anonymous):

Google it if you have to, then find 1/5 of it by multiplying 1/5 by however many pounds it is.

OpenStudy (anonymous):

But the answer is 655.36 lbs. I neec to use a geometric sequence for it

OpenStudy (anonymous):

whaaat, alright im out. I only know how to approximate it. hopefully someone more helpful will come along :P

OpenStudy (anonymous):

Kay, well thanks anyway

OpenStudy (dumbcow):

every hour it loses 1/5 ... so it keeps 4/5 each hour , the geometric sequence is \[(\frac{4}{5})^t\] when t=5 \[= (\frac{4}{5})^{5} = \frac{1,028}{3,125} = 0.32768\] after 5 hours what is left is 0.32768 tons now just convert to pounds

OpenStudy (anonymous):

Ohhh, okay. Thanks

OpenStudy (kropot72):

@wright16 "But the answer is 655.36 lbs". You must have posted an incorrect answer!

OpenStudy (kropot72):

I get the same result as @dumbcow when I solve using the formula for exponential decay.

OpenStudy (anonymous):

Yeah but take that and convert it to lbs. (Multiply by 2000) and you get 655.36

OpenStudy (kropot72):

Where did you get that conversion 1 ton = 2000 lbs ?

OpenStudy (anonymous):

I already knew a ton is 2000lbs

OpenStudy (kropot72):

You are using the conversion for the Short Ton (US).

OpenStudy (anonymous):

Yes